The Department of Agriculture of Economica (DAE) administers the floor price of milk at $4 per pound of milk. To support the price of milk at the price floor, the DAE had to buy up the surplus.

Suppose the Economica’s market demand and supply of milk are as given below:

Q = 120 – 20P (Market demand)

Q = 20P (Market supply)

Refer to the above-market demand and supply equations to answer the following questions.

i) Use the equations to draw the market demand and supply curves and determine the

equilibrium quantity and equilibrium for milk. 


1
Expert's answer
2022-01-02T18:19:53-0500

Put; P = 1, 2, 3, 4 or 5

Market demand:

"Q=120-20P"

"Q_d=120-(20\times1)=100"

"Q_d=120-(20\times2)=80"

"Q_d=120-(20\times3)=60"

"Q_d=120-(20\times4)=40"

"Q_d=120-(20\times5)=20"

Market supply:

"Q=20P"

"Q_s=20\times1=20"

"Q_s=20\times2=40"

"Q_s=20\times3=60"

"Q_s=20\times4=80"

"Q_s=20\times5=100"

Following table market demand and supply against prices:



The market demand and supply curves are shown in graph below:




As shown in above graph, the market demand and supply curves interact with each other when price is $3 per unit. This is called market equilibrium when demand equals to supply. As per above graph, equilibrium quantity for Milk is 60 and equilibrium price is $3.